OpenStudy (anonymous):

how long does it take light to reach us from the sun, 1.49 X 10 to the 8th km away

OpenStudy (anonymous):

a red light has a wavelenght of 620nm. what is the frequency?

OpenStudy (anonymous):

you can use v=s/t for the first and \(v=f \lambda\) for the second.

OpenStudy (anonymous):

note: light has a speed of \(3 \times 10^8m/s\)

OpenStudy (unklerhaukus):

\[v=\frac dt\quad \implies\quad t=\frac dv\] \[d=1.49\times10^8 [\text {km}]=1.49\times10^{11}[\text m]\]\[v=c=3\times10^8 [\text{m/s}]\]

OpenStudy (anonymous):

Thank you!